OpenStudy (jazzyfa30):

Write your answer in the formx=2:3x-9=5(x-3) how do I solve this?

OpenStudy (johnweldon1993):

So I believe it wants you to solve for x here 3x - 9 = 5(x - 3) first...distribute the 5 into the parenthesis...what do you have after that?

OpenStudy (jazzyfa30):

add 15 to both sides

OpenStudy (johnweldon1993):

Right yeah you can do that then we'll have 3x - 9 = 5x - 15 +15 +15 3x + 6 = 5x now what?

OpenStudy (jazzyfa30):

subtract 3x from both side

OpenStudy (johnweldon1993):

right 2x = 6 and finally x = ...?

OpenStudy (jazzyfa30):

x+3

OpenStudy (johnweldon1993):

I'll assume that should be x = 3 ?

OpenStudy (jazzyfa30):

yes sorry

OpenStudy (johnweldon1993):

no problem...and that is your 'x'

OpenStudy (jazzyfa30):

thank you I got the answer right

OpenStudy (johnweldon1993):

perfect...good job!

OpenStudy (jazzyfa30):

I got another one for you

OpenStudy (jazzyfa30):

Josh and Pam have purchased an older home in need of repairs. They have budgeted $3500 for repairs. After spending $1000 for furnace repairs they need to replace 5 windows. What is the maximum they can spend on each of the five windows? Write your answer in the form of w=$300.

OpenStudy (johnweldon1993):

So we should set this up they have 3500 total for repairs. already spent 1000 so they have 2500 left they have 5 windows to replace....how much can each window cost 5w = 2500 how do you solve for w...?